OpenStudy (anonymous):

Find the price per unit (p) that produces the maximum profit P, if the cost function is C=4000 - 40x + .02x^2 and the demand function is p = 50-x/100

OpenStudy (helder_edwin):

ok. u have \[ \large C(x)=4000-40x+.02x^2 \] and \[ \large p=50-\frac{x}{100} \] how do u define profit?

OpenStudy (anonymous):

\[P=x \times p -c\]

OpenStudy (anonymous):

\[P=x(500 -\frac{ x }{ 100}) - *4000 -40x + .02x ^{2}\]

OpenStudy (helder_edwin):

profit is income minus cost. right?

OpenStudy (anonymous):

yeahwhat i posted up there :)

OpenStudy (helder_edwin):

sorry. i told u.

OpenStudy (helder_edwin):

is it 500 or 50 in the priec equation.

OpenStudy (anonymous):

50

OpenStudy (helder_edwin):

ok so we have \[ \large P(x)=px-C(x)=x(50-x/100)-(4000-40x+.02x^2) \]

OpenStudy (anonymous):

yeah i got that :)

OpenStudy (helder_edwin):

\[ \large P(x)=50x-x^2/100-4000+40x-.02x^2 \] \[ \large P(x)=-.03x^2+90x-4000 \] right?

OpenStudy (anonymous):

yeah

OpenStudy (helder_edwin):

great now find the first derivative

OpenStudy (helder_edwin):

and solve 0=P'(x)

OpenStudy (anonymous):

p=50-1500/100 p=$35

OpenStudy (helder_edwin):

everything is correct. sorry. the line crashed.

OpenStudy (helder_edwin):

x=1500 p=35

OpenStudy (anonymous):

hahah its alrite :P

OpenStudy (anonymous):

can u help me with another one? ^^